find the slope if it exists (-11, -8) and (-12,-10)

Answer by mananth(12270)

(-11, -8) and (-12,-10)

..
slope m = (y2-y1)/(x2-x1)
m=-10-(-8)/-12-(-11)
m=-10+8/-12+11
m=-2/-1
m=2